OpenStudy (anonymous):

Any help would be greatly appreciated if f(x)= x^2-x+2 then find f(x+h)-f(x)/h

jimthompson5910 (jim_thompson5910):

\[\Large f(x)= x^2-x+2\] \[\Large f(x+h)= (x+h)^2-(x+h)+2\] \[\Large f(x+h)= x^2+2xh+h^2-x-h+2\]

jimthompson5910 (jim_thompson5910):

\[\Large \frac{f(x+h)-f(x)}{h}\] \[\Large \frac{(x^2+2xh+h^2-x-h+2)-(x^2-x+2)}{h}\] I'll let you simplify that.

OpenStudy (anonymous):

Sorry, I walked away from the computer for a few, thanks so much!! I appreciate the explanation, this way I should be able to figure out the rest as well!!

jimthompson5910 (jim_thompson5910):

you're welcome